OpenStudy (cinnamonbun):

The prism has a volume of 120 cm3. What would the volume of the prism be if each dimension were halved? 60?

OpenStudy (cinnamonbun):

@TheSmartOne

jhonyy9 (jhonyy9):

volume of prism = area of base time height soppose a prism with base a square and height equal length of square side 120 = a*a*a so than eache dimensions was halved 120 = (a/2)*(a/2)*(a/2) 120 = (a/2)^3 120 = a^3 /8 a^3 = 120/8 = 15 a= cuberoot(15) so than a side length and the height will be equal cuberoot 15 so than the volume will be (cuberoot15)^3 = 15 cm^3 hope this is understandably easy good luck bye bye